# Classical Combinatorial Optimization Scaling for Random Ising Models on 2D Heavy-Hex Graphs

### v1

@@ Abstract
Motivated by near term quantum computing hardware limitations, combinatorial optimization problems that can be addressed by current quantum algorithms and noisy hardware with little or no overhead are used to probe capabilities of quantum algorithms such as the Quantum Approximate Optimization Algorithm (QAOA). In this study, a specific class of near term quantum computing hardware defined combinatorial optimization problems, Ising models on heavy-hex graphs both with and without geometrically local cubic terms, are examined for their classical computational hardness via empirical computation time scaling quantification. Specifically the Time-to-Solution metric using the classical heuristic simulated annealing is measured for finding optimal variable assignments (ground states), as well as the time required for the optimization software Gurobi to find an optimal variable assignment. Because of the sparsity of these Ising models, the classical algorithms are able to find optimal solutions efficiently even for large instances (i.e. $100,000$ variables). The Ising models both with and without geometrically local cubic terms exhibit average-case linear-time or weakly quadratic scaling when solved exactly using Gurobi, and the Ising models with no cubic terms show evidence of exponential-time Time-to-Solution scaling when sampled using simulated annealing. These findings point to the necessity of developing and testing more complex, namely more densely connected, optimization problems in order for quantum computing to ever have a practical advantage over classical computing. Our results are another illustration that different classical algorithms can indeed have exponentially different running times, thus making the identification of the best practical classical technique important in any quantum computing vs. classical computing comparison.
1 Introduction
@@
The Quantum Approximate Optimization Algorithm (QAOA) is a quantum heuristic for solving combinatorial optimization problems [1, 2]. QAOA is a promising quantum algorithm as it can be executed on noisy current quantum computers that are not error-corrected and some signal can still be measured despite the noise. There is evidence that QAOA could be competitive with classical algorithms [3, 4, 5, 6, 7, 8, 9, 10]. QAOA and similar quantum algorithms that solve combinatorial optimization problems are of considerable interest because combinatorial optimization problems are a fundamentally challenging and important computational task in information processing – and therefore if quantum mechanical effects can be utilized to efficiently solve such problems, then this would be an excellent use for quantum computing. However, studying quantum algorithms performance such as QAOA is inherently challenging because simulations of the algorithm requires either classical simulation methods, which are severely limited either due to the exponential classical computational complexity of simulating many-body quantum systems, or requires using current state of the art quantum computers which are heavily affected by limited qubit coherence times and various sources of control error. Thus, generally simulations to study QAOA performance are limited to either i) small problem sizes, or ii) intermediate-scale hardware with noise, or both. This limitation of studying quantum algorithms has motivated several different studies, including may different tasks not necessarily only for combinatorial optimization, which define problem types that can be implemented on NISQ (Noisy Intermediate-Scale Quantum) [11] hardware with little or no overhead with respect to compute time or qubit use [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. These types of studies allow various quantum algorithms to be investigated in a regime where some amount of signal can be obtained from the noisy quantum hardware, but the problem instances are not very computationally hard or particularly computationally relevant. In this study, one of these types of NISQ-friendly problem instances is examined for its computational hardness using classical algorithms using the commercial optimization software Gurobi [26] and the general-purpose heuristic simulated annealing [27]. The problem instances are a type of discrete optimization problem that are random Ising models, where the goal is to find the global minimum energy of the Ising model or equivalently an optimal decision variable assignment.
@@
Specifically, this study examines the classical computational hardness of ground-state finding for two classes of random spin glass Ising models defined on heavy-hex graphs, those with only linear and quadratic polynomial terms and those with additional cubic (three body) terms, were introduced and defined in refs. [14, 15]. This class of Ising models is designed to be extremely hardware compatible for current IBM Quantum superconducting qubit processors which have two-qubit gate operations that define a sparse graph structure that is generally known as heavy-hex [28]. These heavy-hex Ising models, and in particular the implementation of the short-depth QAOA circuits that can sample these Ising models, have been subsequently been used in several studies [29, 16, 30, 31].
2 Methods
@@
First, we created a custom generator that produces heavy-hex graphs – the structure of these graphs is a tileable lattice [28]. This allows us to create graphs with an arbitrary number of nodes. We focus specifically on 2-dimensional heavy-hex structures that are approximately square (approximately the same number of hexagonal unit cells on all sides of the 2D lattice) because this follows the existing structures of current IBM quantum processors [28]. Figure 1 shows three of these generated heavy-hex graphs. Based on these graph structures, next the Ising models defined on those graphs can be created. The heavy-hex Ising model generator uses the Networkx package [32].
The heavy-hex graph spin glass Ising models were generated using the same methods outlined in refs. [14, 15, 16], but applied to significantly larger problem sizes. These Ising models are minimization and unconstrained combinatorial optimization problems.
@@
Typically we will refer to the cost value for a given sample as the energy, where the goal is to minimize the energy. Every heavy-hex graph is a bipartite graph that has vertices $V=\{1,\ldots,n\}$ and can be uniquely bipartitioned as $V=V_{2}\sqcup V_{3}$ with $E\subset V_{2}\times V_{3}$, where $V_{i}$ has vertices of greatest degree $i$. $W$ is the set of vertices $l$ in $V_{2}$ that have degree $2$ and the neighbors are $n_{1}(l)$ and $n_{2}(l)$. $d_{v}$, $d_{i,j}$, and $d_{l,n_{1}(l),n_{2}(l)}$ are the linear, quadratic and cubic coefficients. The Ising model coefficients $d$ are randomly drawn from $\{+1,-1\}$ with probability $0.5$. In this study we examine the computational hardness of ground-state finding for both the Ising models with and without the geometrically local cubic terms.
The primary classical solver that will be used in this study is the commercial Gurobi optimization software [26] - which is not a specific algorithm, but rather a piece of software that uses a variety of techniques [33, 34, 35, 36, 37], such as branch-and-bound, and is designed to deterministically find an optimal solution, with a provable optimality gap so that it can be guaranteed there exists no better variable assignment (given sufficient compute time). Gurobi is not restricted to solving only discrete optimization problems, but here we use it to solve these discrete optimization problems where the variable states or either $-1$ or $+1$.
Gurobi does not directly handle higher order terms in the objective function polynomial, and therefore they must be structured so as to be compatible with the software. We utilize two different formulation techniques to address the cubic terms and obtain global optimal solutions of the optimization problems:
@@
    The cubic $ZZZ$ terms in the higher order Ising models are order-reduced by introducing auxiliary variables, and quadratic interactions with those auxiliary variables, that preserve the optimal solution(s) to the original problem. This order reduction is performed using the heuristic order reduction method in the Python 3 module dimod [^1] https://github.com/dwavesystems/dimod. This order reduction is deterministic meaning that the resulting linear and quadratic terms in the polynomial are always the same given the same set of parameters (and there is no random seed). There is a penalty weight parameter that must be tuned such that the optimal (globally minimum) solutions are preserved in the order-reduced problem. Setting this penalty weight optimally is computationally hard in general. In this case, we set the weights to a large positive value equal to the number of nodes in the heavy-hex graph since we have found that in general as long as the weight is sufficiently greater than the largest magnitude coefficient in the original polynomial, the optimal solutions are preserved. This order reduced polynomial contains only linear and quadratic terms - and in Gurobi is formulated as a binary Integer Quadratic Program (IQP), where the variable spins are converted to binary in the cost function as $12z_{i}-$ 1. Importantly, this technique introduces a large overhead of auxiliary variables, and is not very optimized.
@@
    Formulate the polynomial as a binary Integer Linear Program, where the variable states of the optimization problem are first converted into binary states. The higher order terms are lifted [36, 38, 39, 40, 41] into valid inequality constraint representations that do not use the higher order variable interactions directly, by using lower dimensional variable inequalities. In order to convert variable states from spins into binary, with the higher order terms, the Python 3 module dimod is also used, which introduces additional higher order terms. All higher order terms can then be represented as a linear program. This ILP formulation follows the same basic methods used in refs. [12, 42].
The energy of the solution vectors produced by Gurobi when these two different methods are used are compared for all problem instances to ensure that the methods agree. In practice we found that these methods always agreed - however it is not clear whether the penalty weight used for the first method will always ensure an optimal solution is correctly encoded in the order-reduced polynomial.
@@
Figure 3: Gurobi runtime scaling (log scale y-axis) as a function of Ising model problem size (x-axis), where the instance contain no higher order terms. For each system size, $10$ random optimization problems are solved and the runtimes are plotted as individual points; the mean time of each set of $10$ instances is plotted as a connected line.
In addition to being solved exactly using Gurobi, these Ising models are also sampled using an existing C++ implementation of simulated annealing that is available as a Python 3 library dwave-neal [^2] https://github.com/dwavesystems/dwave-neal. The simulated annealing algorithm performs a fixed set of Metropolis-Hastings spin updates, which we vary to be either $100$ or $1000$ where each update uses increasing $\beta$ values. The default geometric simulated annealing schedule is used, which defines the values of $\beta$ at each step of the schedule. $\beta$ is the inverse temperature of the Boltzmann distribution $\frac{1}{k_{B}T}$, $T$ is the thermodynamic temperature, and $k_{B}$ is the Boltzmann constant [27]. The total number of simulated annealing simulations performed for each Ising model, where each simulation generates a single low-energy sample of the optimization problem, is fixed to be $50,000$. The computation quality of the simulated annealing sampling is measured using Time-to-solution (TTS) [43]:
$TTS=\frac{T_{\text{CPU}}}{N_{s}}\frac{\log(1-0.99)}{\log(1-p)},$ (2)
where $p$ denotes the measured optimal solution sampling success rate. $N_{s}$ is the total number of samples obtained, $T_{\text{CPU}}$ is the classical CPU time required to generate the full set of $N_{s}$ samples measured as Python single-process time. TTS was defined in the context of quantifying heuristic sampling quantum and classical algorithm runtimes, and serves as a good runtime complexity measure in the absence of guarantees of optimality [43, 44, 45]. In this case TTS is defined as the expected compute time required to sample an optimal solution of the optimization problem at least once, with $99\%$ confidence. The optimal solution sampling success rate $p$ of finding the ground-state energy in this case is measured using the exact solutions obtained from Gurobi for the same problem instances.
@@
Gurobi Optimizer [26] version 11.0.3 is used for all Gurobi computations. The compute hardware used for simulated annealing simulations as well as Gurobi computations is a Red Hat Linux node with Intel(R) Xeon(R) CPU E5-2695 v4 2.10GHz. All Gurobi simulations use a time limit of $100,000$ seconds, a single thread, and a MIP gap of $1\mathrm{e}{-8}$. All binary solutions found by Gurobi are adapted to spins using the mapping of $1\mapsto 1$, $0\mapsto-1$. Gurobi can be used to find multiple feasible optimal solutions, if they exist, however in this case we use Gurobi to only find a single optimal solution for each problem instance.
Figure 4: Gurobi runtime scaling (log scale y-axis) as a function of Ising model problem size (x-axis) for the instances with geometrically local cubic terms. Left plot shows the runtime scaling for the binary ILP method, and the right plot shows the runtime scaling for the binary IQP order reduction method. These two methods of handling the cubic terms are applied to exactly the same problem instance. At each system size N, $10$ instances are solved by each method, and the individual runtimes are shown as individual points.
Figure 5: Simulated annealing TTS scaling (log scale y-axis) as a function of Ising model problem size (x-axis) for the instances containing no higher order terms. $100$ Metropolis-Hastings updates (left) and $1000$ Metropolis-Hastings updates (right). At each system size N, $10$ different random-coefficient instances are solved and then the average time to solve those $10$ instances are shown as a function of $N$ by the connected green line.
3 Results
Problem type Simulation Function fit type Fitted coefficients RMSE
Quadratic SA TTS, $1000$ updates log $a=22.8,b=0.0001$ $572.6$
Quadratic SA TTS, $1000$ updates linear $a=0.111,b=0.0001$ $510.24$
Quadratic SA TTS, $1000$ updates quadratic $a=0.000032,b=0.000001,c=0.000001$ $464.2$
Quadratic SA TTS, $1000$ updates exponential $a=0.0038,b=0.01,c=0.01$ $172.7$
Quadratic SA TTS, $100$ updates log $a=2.78,b=0.0001$ $55.3$
Quadratic SA TTS, $100$ updates linear $a=0.0385,b=0.0001$ $48.7$
Quadratic SA TTS, $100$ updates quadratic $a=0.000036,b=0.000001,c=0.000001$ $41.8$
Quadratic SA TTS, $100$ updates exponential $a=0.011,b=0.001,c=0.0001$ $6.29$
Quadratic Gurobi IQP log $a=0.158,b=0.0001$ $1.761$
Quadratic Gurobi IQP linear $a=0.0001,b=0.0001$ $0.683$
Quadratic Gurobi IQP quadratic $a=0.000001,b=0.000001,c=0.000001$ $3444.4$
Quadratic Gurobi IQP exponential $a=0.0001,b=0.001,c=1.788$ $2.356$
Cubic Gurobi ILP log $a=102.597,b=-1186.39$ $230.7$
Cubic Gurobi ILP linear $a=0.00791,b=0.0001$ $27.58$
Cubic Gurobi ILP quadratic $a=0.000001,b=0.000001,c=0.000001$ $2940.4$
Cubic Gurobi ILP exponential $a=0.0001,b=1.1,c=122.3$ $153.60$
Cubic Gurobi IQP log $a=147.3,b=0.0001$ $2040.8$
Cubic Gurobi IQP linear $a=0.12499,b=0.0001$ $315.67$
Cubic Gurobi IQP quadratic $a=0.000001,b=0.0802,c=0.000001$ $192.6$
Cubic Gurobi IQP exponential $a=0.0001,b=140,c=691$ $947.1$
@@ 3 Results
Table 1: Curve fitting for the computation time with respect to system size for the different optimization problem types and solvers. Linear function is fit to $ax+b$, log function is fit to $a\log_{2}(x)+b$, quadratic function is fit to $ax^{2}+b\cdot x+c$, and exponential function is fit to $2^{xa}b+c$. Where $x$ denotes the system size and the coefficients $a$, $b$, $c$ are fit using least squares optimization. The dependent variable which these curves are being fit to, for each system size $x$, is the raw compute time in the case of Gurobi computations and is the TTS in the case of simulated annealing. Bounds are used to restrict the least squares fitting to not set the fitting variables too close to 00 with a minimum coefficient precision limit of $1\cdot 10^{-6}$, for example in the exponential and quadratic cases, which would effectively be replicating a linear function scaling for small system sizes. All fitting coefficients are also forced to be non-negative. The lowest RMSE function fit is highlighted as red text for each problem type and solver setting.

### v2

@@ Abstract
Motivated by near term quantum computing hardware limitations, combinatorial optimization problems that can be addressed by current quantum algorithms and noisy hardware with little or no overhead are used to probe capabilities of quantum algorithms such as the Quantum Approximate Optimization Algorithm (QAOA). In this study, a specific class of near term quantum computing hardware defined combinatorial optimization problems, Ising models on heavy-hex graphs both with and without geometrically local cubic terms, are examined for their classical computational hardness via empirical computation time scaling quantification. Specifically the Time-to-Solution metric using the classical heuristic simulated annealing is measured for finding optimal variable assignments (ground states), as well as the time required for the optimization software Gurobi to find an optimal variable assignment. Because of the sparsity of these Ising models, the classical algorithms are able to find optimal solutions efficiently even for large instances (i.e., $100,000$ spin variables). The Ising models both with and without geometrically local cubic terms exhibit average-case linear-time or weakly quadratic scaling when solved exactly using Gurobi, and the Ising models with no cubic terms show evidence of exponential-time Time-to-Solution scaling when sampled using simulated annealing. These findings point to the necessity of developing and testing more complex, namely more densely connected, optimization problems in order for quantum computing to ever have a practical advantage over classical computing. Our results are another illustration that different classical algorithms can indeed have exponentially different running times, thus making the identification of the best practical classical technique important in any quantum computing vs. classical computing comparison.
I Introduction
@@
The Quantum Approximate Optimization Algorithm (QAOA) is a quantum heuristic for solving combinatorial optimization problems [1, 2]. QAOA is a promising quantum algorithm as it can be executed on noisy current quantum computers that are not error-corrected and some signal can still be measured despite the noise. There is evidence that QAOA could be competitive with classical algorithms [3, 4, 5, 6, 7, 8, 9, 10]. QAOA and similar quantum algorithms that solve combinatorial optimization problems are of considerable interest because combinatorial optimization problems are a fundamentally challenging and important computational task in many fields including information processing – and therefore if quantum mechanical effects can be utilized to efficiently solve such problems, then this would be an excellent use for quantum computing. However, studying quantum algorithms performance such as QAOA is inherently challenging because simulations of the algorithm requires either classical simulation methods, which are severely limited either due to the exponential classical computational complexity of simulating many-body quantum systems, or requires using current state of the art quantum computers which are heavily affected by limited qubit coherence times and various sources of control error. Thus, generally simulations to study QAOA performance are limited to either i) small problem sizes, or ii) intermediate-scale hardware with noise, or both. This limitation of studying quantum algorithms has motivated several different studies, including many different computational tasks not necessarily only for combinatorial optimization, which define problem types that can be implemented on NISQ (Noisy Intermediate-Scale Quantum) [11] hardware with little or no overhead with respect to compute time or qubit use [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. These types of studies allow various quantum algorithms to be investigated in a regime where some amount of signal can be obtained from the noisy quantum hardware, but these problem instances are typically not very computationally hard (with a few exceptions, such as ref. [12]) or particularly computationally relevant to a specific industry combinatorial optimization problem.
In this study, one of these types of NISQ-friendly problem instances is examined for its computational hardness using classical algorithms. We focus on two approaches: the commercial optimization software Gurobi [26] and the heuristic Simulated Annealing [27]. Simulated annealing is very general-purpose, meaning it can work quite well on a variety of different types of optimization problems, and therefore is typically considered as a performance baseline. Gurobi on the other hand serves as a powerful mathematical optimization solver tool, which in general we expect is closer to being state of the art for general purpose optimization. We note that simulated annealing is not a state of the art heuristic solver of optimization problems – there exist algorithms which perform better than simulated annealing (especially when considering specific types of optimization problems), such as Simulated Quantum Annealing [28], Tabu Search [29], Parallel Tempering [30] and others, which for example are benchmarked in ref. [12].
@@
The problem instances that we consider in this study are a type of discrete optimization problem, specifically random Ising models. The computational goal is to find the global minimum energy of the Ising model, or equivalently an optimal decision variable assignment. Specifically, this study examines the classical computational hardness of ground-state finding for two classes of random spin glass Ising models defined on heavy-hex graphs, those with only linear and quadratic polynomial terms and those with additional cubic (three body) terms, were introduced and defined in refs. [14, 15]. This class of Ising models is designed to be extremely hardware compatible for current IBM Quantum superconducting qubit processors which have two-qubit gate operations that define a sparse graph structure that is generally known as heavy-hex [31, 32, 33]. These heavy-hex Ising models, and in particular the implementation of the short-depth QAOA circuits that can sample these Ising models, have been subsequently been used in several studies [34, 16, 35, 36].
II Methods
@@
First, we created a custom generator that produces heavy-hex graphs – the structure of these graphs is a tileable lattice [31]. This allows us to create graphs with an arbitrary number of nodes. We focus specifically on 2-dimensional heavy-hex structures that are approximately square (approximately the same number of hexagonal unit cells on all sides of the 2D lattice) because this follows the existing structures of current IBM quantum processors [31]. Figure 1 shows three of these generated heavy-hex graphs. Based on these graph structures the Ising models defined on those graphs can be created. The heavy-hex Ising model generator uses the Networkx package [37].
The heavy-hex graph spin glass Ising models were generated using the same methods outlined in refs. [14, 15, 16], but applied to significantly larger problem sizes. In the context of the field of combinatorial optimization, these Ising models are can be classified as minimization and unconstrained combinatorial optimization problems (as opposed to, for example, a type of constrained optimization problem).
@@
Typically we will refer to the cost value for a given sample as the energy, where the goal is to minimize the energy. Every heavy-hex graph is a bipartite graph that has vertices $V=\{1,\ldots,n\}$ and can be uniquely bipartitioned as $V=V_{2}\sqcup V_{3}$ with $E\subset V_{2}\times V_{3}$, where $V_{i}$ has vertices of greatest degree $i$. $W$ is the set of vertices $l$ in $V_{2}$ that have degree $2$ and the neighbors are $n_{1}(l)$ and $n_{2}(l)$. $d_{v}$, $d_{i,j}$, and $d_{l,n_{1}(l),n_{2}(l)}$ are the linear, quadratic and cubic coefficients. The Ising model coefficients $d$ are randomly drawn from $\{+1,-1\}$ with probability $0.5$. This type of Ising model could be considered a type of sparse ±$\pm J$ model [38, 39]. In this study we examine the computational hardness of ground-state finding for the Ising models with and without the geometrically local cubic terms.
The primary classical solver that will be used in this study is the commercial Gurobi optimization software [26] - which is not a single algorithm, but rather a piece of software that uses a variety of techniques [40, 41, 42, 43, 44], such as branch-and-bound, and is designed to deterministically find an optimal solution, with a provable optimality gap so that it can be guaranteed there exists no better variable assignment (given sufficient compute time). Gurobi is not restricted to solving only discrete optimization problems, but here we use it to solve these discrete optimization problems where the variable states or either $-1$ or $+1$.
Gurobi does not directly handle higher order terms in the objective function polynomial where the variable states are strictly spins ${+1,-1}$, and therefore they must be structured so as to be compatible with the software. We utilize two different formulation techniques to address the cubic terms and obtain global optimal solutions of the optimization problems:
@@
    The cubic $ZZZ$ terms in the higher order Ising models are order-reduced by introducing auxiliary variables, and quadratic interactions with those auxiliary variables, that preserve the optimal solution(s) to the original problem. This order reduction is performed using the heuristic order reduction method in the Python 3 module dimod [^1] https://github.com/dwavesystems/dimod. This order reduction is deterministic meaning that the resulting linear and quadratic terms in the polynomial are always the same given the same set of parameters (and there is no random seed). There is a penalty weight parameter that must be tuned such that the optimal (globally minimum) solutions are preserved in the order-reduced problem. Setting this penalty weight optimally is computationally hard in general. In this case, we set the weights to a large positive value equal to the number of nodes in the heavy-hex graph since we have found that in general as long as the weight is sufficiently greater than the largest magnitude coefficient in the original polynomial, the optimal solutions are preserved. There are very likely alternate order reduction techniques which do not require using a large penalty weight, however, order reduction methods are not the specific focus of this study and in particular this method empirically guarantees optimality in the tests we performed. This order reduced polynomial contains only linear and quadratic terms - and in Gurobi is formulated as a binary Integer Quadratic Program (IQP), where the binary variables (which are the variable types that Gurobi handles) are converted into spins via $12x_{i}-$ 1. Importantly, this technique introduces a large overhead of auxiliary variables, and in particular is not heavily optimized. More complex and tailored order reduction could certainly be applied to these types of higher order terms, but here we utilize general purpose heuristics that work efficiently.
@@
    Formulate the polynomial as a binary Integer Linear Program (ILP), where the variable states of the optimization problem are first converted into binary states. In order to convert variable states from spins into binary, with the higher order terms, the Python 3 module dimod is also used, which introduces additional higher order terms. The higher order terms are then linearized [46, 47, 48, 49] into valid inequality constraint representations that do not use the higher order variable interactions directly, by using lower dimensional variable inequalities. An example is given below. All higher order terms can then be represented as a binary Integer Linear Program. This ILP formulation follows the same basic methods used in refs. [12, 50].
Here we show a small example of linearization of a cubic term, which is used in the binary ILP formulation of the problem instances. If we want to linearize the following cubic term,
$K=x_{i}x_{j}x_{k},$
where $x_{i},x_{j},x_{k}$ are binary variables. Then we can specify a logically equivalent formulation using the following linear constraints:
$K$ $\leq x_{i}$
$K$ $\leq x_{j}$
$K$ $\leq x_{k}$
$K$ $\geq x_{i}+x_{j}+x_{k}-2$
The energy of the solution vectors produced by Gurobi when these two different methods are used, are compared against each other (for a given problem instance) to ensure that the two methods agree. In practice we found that these methods always agreed - however it is not clear whether the penalty weight used for the first method will always ensure an optimal solution is correctly encoded in the order-reduced polynomial for arbitrary polynomials.
In addition to being solved exactly using Gurobi, these Ising models are also sampled using an existing C++ implementation of simulated annealing that is available as a Python 3 library dwave-neal [^2] https://github.com/dwavesystems/dwave-neal. The simulated annealing algorithm performs a fixed set of Metropolis-Hastings spin updates, which we vary to be either $100$ or $1000$ where each update uses a schedule of increasing $\beta$ values. The default geometric simulated annealing schedule is used, which defines the values of $\beta$ at each step of the schedule. $\beta$ is the inverse temperature of the Boltzmann distribution $\frac{1}{k_{B}T}$, $T$ is the thermodynamic temperature, and $k_{B}$ is the Boltzmann constant [27]. The total number of simulated annealing simulations performed for each Ising model, where each simulation generates a single low-energy sample of the optimization problem, is fixed to be $50,000$. The goal was to use a fixed number of samples that was reasonably large so that finite sampling does not dominate the reported statistics, but also computationally feasible to carry out the simulations in a reasonable amount of time. We use simulated annealing sampling strictly on the Ising models with no cubic terms, so as to avoid the increased computational overhead that comes from order reduction of the cubic terms (this particular simulated annealing solver works only with linear and quadratic terms). The computation quality of the simulated annealing sampling is measured using the well established heuristic probabilistic sampler based compute time metric Time-to-solution (TTS) [52]:
$TTS=\frac{T_{\text{CPU}}}{N_{s}}\frac{\log(1-0.99)}{\log(1-p)},$ (2)
where $p$ denotes the measured optimal solution sampling success rate (sampling rate of the optimal cost value, not any particular variable configuration). The optimal solution cost value is found using Gurobi. $N_{s}$ is the total number of samples obtained, $T_{\text{CPU}}$ is the classical CPU time required to generate the full set of $N_{s}$ samples measured as Python single-process time. TTS was defined in the context of quantifying heuristic probabilistic sampling quantum and classical algorithm runtimes, and serves as a good runtime complexity measure in the absence of guarantees of optimality [52, 53, 54]. In this case TTS is defined as the expected compute time required to sample an optimal solution of the optimization problem at least once, with $99\%$ confidence. The optimal solution sampling success rate $p$ of finding the ground-state energy in this case is measured using the exact solutions obtained from Gurobi for the same problem instances.
@@
Gurobi Optimizer [26] version 11.0.3 is used for all Gurobi computations. The compute hardware used for simulated annealing simulations as well as Gurobi computations is a Red Hat Linux node with Intel(R) Xeon(R) CPU E5-2695 v4 2.10GHz. All Gurobi simulations use a time limit of $100,000$ seconds, a single thread, and a MIP gap of $1\mathrm{e}{-8}$. Gurobi was run with the intention of converging on an optimal solution, and each Gurobi run was checked for this condition where the optimality gap had closed and therefore the solution that we have is a verified optimal solution. At some point, for sufficiently large problems, the Gurobi simulation time would exceed this time limit that was specified, but for all simulations reported on in this study, each individual run completed with a converged optimality gap within the specific time limit. All binary solutions found by Gurobi are adapted to spins using the mapping of $1\mapsto 1$, $0\mapsto-1$. Gurobi can be used to find multiple feasible optimal solutions, if they exist, however in this case we use Gurobi to only find a single optimal solution for each problem instance. We use only a single thread in order to make the compute time scaling measurements very clear, and in particular to avoid multi-threaded CPU time measurements.
@@
All curve fits are performed using non linear least-squares fitting using the scipy [55] Python 3 library. Bounds are used to restrict the least squares fitting to not set the fitting variables too close to 00 with a minimum coefficient precision limit of $10^{-7}$, for example in the exponential and quadratic cases, which would effectively be replicating a linear function scaling for small system sizes. All fitting coefficients are also forced to be non-negative. The linear function is fit to $y=ax+b$, log function is fit to $y=a\ln{x}+b$, quadratic function is fit to $y=ax^{2}+b\cdot x+c$, and exponential function is fit to $y=e^{xa}\cdot b$.
Figure 3: Gurobi runtime scaling (log scale y-axis) as a function of Ising model problem size (x-axis), where the instance contain no higher order terms. The x-axis (problem size) is exactly the number of nodes in the heavy-hex graph. For each system size, $10$ random optimization problems are solved and the runtimes are plotted as individual points; the mean time of each set of $10$ instances is plotted as a connected line. The problem formulation used in these simulations was binary IQP.
Figure 4: Gurobi runtime scaling (log scale y-axis) as a function of Ising model problem size (x-axis) for the instances with geometrically local cubic terms. The x-axis (problem size) is exactly the number of nodes in the heavy-hex graph. Left plot shows the runtime scaling for the binary ILP method, and the right plot shows the runtime scaling for the binary IQP order reduction method. These two methods of handling the cubic terms are applied to exactly the same problem instance. At each system size N, $10$ instances are solved by each method, and the individual runtimes are shown as individual points. Notice that the binary IQP method (right) requires approximately an order of magnitude more compute time than the binary ILP method (left) at $60,000$ variable problem sizes.
III Results
Problem type Simulation Function type Function fit Precise coefficient fits RMSE
Quadratic SA TTS, $1000$ updates log $37.2\,\ln x$ $a=37.2,\;b=10^{-6}$ 608.9
linear $0.123\,x$ $a=0.123,\;b=10^{-7}$ 534.3
quadratic $3.2\cdot 10^{-5}\,x^{2}$ $a=3.2\cdot 10^{-5},\;b=c=10^{-7}$ 464.2
exponential $0.022\,e^{0.002\,x}$ $a=0.002,\;b=0.022$ 235.7
SA TTS, $100$ updates log $4.01\,\ln x$ $a=4.01,\;b=10^{-6}$ 55.3
linear $0.0385\,x$ $a=0.0385,\;b=10^{-7}$ 48.7
quadratic $3.6\cdot 10^{-5}\,x^{2}$ $a=3.6\cdot 10^{-5},\;b=c=10^{-7}$ 41.8
exponential $0.000128\;e^{0.00898\,x}$ $a=0.00898,\;b=0.000128$ 4.5
Gurobi IQP log $0.24\,\ln x$ $a=0.24,\;b=10^{-6}$ 2.7
linear $9\!\cdot\!10^{-5}\,x$ $a=9\!\cdot\!10^{-5},\;b=10^{-7}$ 0.62
quadratic 0 $a=b=c=10^{-7}$ 373.1
exponential $0.0001\;e^{0.0001\,x}$ $a=b=0.0001$ 135.5
Cubic Gurobi ILP log $25.39\,\ln x$ $a=25.39,\;b=10^{-6}$ 230.7
linear $0.00791\,x$ $a=0.00791,\;b=10^{-7}$ 27.58
quadratic 33.4 $a=b=10^{-7},\;c=33.4$ 56.7
exponential $0.065\;e^{0.0001\,x}$ $a=0.0001,\;b=0.065$ 230.5
Gurobi IQP log $226.5\,\ln x$ $a=226.5,\;b=10^{-6}$ 2187.6
linear $0.127\,x$ $a=0.127,\;b=10^{-7}$ 329.7
quadratic $7\cdot 10^{-7}\,x^{2}+0.09\,x$ $a=7\cdot 10^{-7},\;b=0.09,\;c=10^{-7}$ 171.9
exponential $19.7\;e^{0.0001\,x}$ $a=0.0001,\;b=19.7$ 1723.2
Table 1: Curve fitting for the computation time with respect to system size for the different optimization problem types and solvers. The log function is $y=a\ln{x}+b$, the linear function is $y=ax+b$, the quadratic function is $y=ax^{2}+b\cdot x+c$, and the exponential function is $y=e^{xa}\cdot b$. Here $x$ denotes the system size (specifically, the number of spins in the underlying Ising model) and the coefficients $a$, $b$, $c$ are fit using least squares optimization. The dependent variable which these curves are being fit to, for each system size $x$, is the raw compute time in the case of Gurobi computations and is the TTS in the case of simulated annealing (average compute time over the $10$ random instances for each problem size). The lowest RMSE function fit is highlighted as red text for each problem type and solver setting.